# Scaling Limit of a Stochastic Clustering Model on $\mathds{R}$

### v1

@@ 3 Proofs of Main Results
and thus $\overleftarrow{L}^{(t)}\to 0$ in $L^{2}$. It follows that $\langle\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)},\overleftarrow{\Gamma}^{(0)}\rangle\to\overleftarrow{M}^{(\infty)}$ in $L^{2}$. This establishes the single-dimensional distribution convergence since the choice of $\overleftarrow{\eta}^{(0)}=e_{0}$ is without loss of generality.
Finally, we extend to finite-dimensional distribution convergence. Indeed, for $\overleftarrow{\eta_{0}}=\sum_{i\in F}e_{i}$, where $F$ is finite, we can construct a finite-dimensional martingale driven by the same renewal processes $\rho^{(t)}$, each of which converges a.s. and in $L^{2}$. This last statement follows from shift-invariance and using the previous argument for $\overleftarrow{\eta}_{0}=e_{0}$. It is easily seen that the expectation of all linear functionals of the finite-dimensional process also converge; this establishes the claimed finite-dimensional convergence.
@@
The proof of this result is almost identical to that of Theorem 2.1. When defining the reverse-time process, instead of using the tuples $(\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$ and $(\overleftarrow{\eta_{i}}^{(t)},2\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$, we instead use the tuples $(\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$ and $(0,2\overleftarrow{\eta_{i}}^{(t)},2\overleftarrow{\eta_{i}}^{(t)})$.
@@
This result follows by checking that this is true when $\eta^{(0)}=e_{i}$ for $i\in\mathds{Z}$ and then appealing to linearity of expectation. Recall that for $h(\cdot,\cdot)$ defined as the inner product in the corollary statement, the Markov process $\Gamma^{(\cdot)}$ and $\eta^{(\cdot)}$ are dual if for all $t,\Gamma^{(0)},\eta^{(0)}$, we have:
$\operatorname{\mathds{E}}_{\Gamma^{(0)}}h(\eta^{(0)},\Gamma^{(t)})=\operatorname{\mathds{E}}_{\eta^{(0)}}h(\eta^{(t)},\Gamma^{(0)}).$
For $\overleftarrow{\eta}^{(0)}=e_{i}$, the result is shown in the proof of Theorem 2.1.
@@
This result follows by taking the proofs of Theorems 2.1 and 2.3, and coupling the point processes used to determine the un-merging indices.
4 A Further Conjecture
Let $\mathcal{G}_{+}$ denote the space of non-negative integer-valued sequences indexed by $\mathds{Z}_{+}$.
Conjecture 1.
@@ 4 A Further Conjecture
Then $\overleftarrow{F}^{(t)}\Rightarrow\overleftarrow{F}^{(\infty)}$, where the (random) mass of $\overleftarrow{F}^{(\infty)}$ is distributed as $\overrightarrow{\Xi}^{(\infty)}_{1}-\overrightarrow{\Xi}^{(\infty)}_{0}$, and the length of the (random) support of $\overleftarrow{F}^{(\infty)}$ is distributed as $\overrightarrow{G}^{(\infty)}_{1}$.
We explicitly write out a dynamics for $\overleftarrow{F}^{(t)}$ in the following lemma.
Lemma 4.1.
We have for all $t\geqslant 0,x\in\mathds{R}$,
$\overleftarrow{F}^{(t)}(x)+\left(\frac{3}{4}\right)^{(t+1)}\sum_{n\leqslant\left(\frac{4}{3}\right)^{t}x}{\left(\frac{4}{3}\right)^{t}\left[\overleftarrow{F}^{(t)}\left(\left(\frac{3}{4}\right)^{t}n\right)-\overleftarrow{F}^{(t)}\left(\left(\frac{3}{4}\right)^{t}(n-1)\right)\right]\left(\mathds{1}_{n\in\rho^{(t+1)}}-\frac{1}{3}\right)}$
$=\overleftarrow{F}^{(t+1)}\left(x+\left(\frac{3}{4}\right)^{t+1}\left(\left|\rho^{(t+1)}\cap\left[0,\left(\frac{4}{3}\right)^{t}x\right]\right|-\frac{1}{3}\left(\frac{4}{3}\right)^{t}x\right)\right).$
For notational simplicity, we denote these dynamics as:
where
$\alpha^{(t+1)}_{x}:=\left(\frac{3}{4}\right)^{(t+1)}\sum_{n\leqslant\left(\frac{4}{3}\right)^{t}x}{\left(\frac{4}{3}\right)^{t}\left[\overleftarrow{F}^{(t)}\left(\left(\frac{3}{4}\right)^{t}n\right)-\overleftarrow{F}^{(t)}\left(\left(\frac{3}{4}\right)^{t}(n-1)\right)\right]\left(\mathds{1}_{n\in\rho^{(t+1)}}-\frac{1}{3}\right)}$
and
$\beta^{(t+1)}_{x}:=\left(\frac{3}{4}\right)^{t+1}\left(\left|\rho^{(t+1)}\cap\left[0,\left(\frac{4}{3}\right)^{t}x\right]\right|-\frac{1}{3}\left(\frac{4}{3}\right)^{t}x\right).$
Proof.
Fix $t$ and consider some fixed integer $n\in\left[0,\left(4/3\right)^{t}\right]$. There is
$\overleftarrow{Z}_{n}^{(t)}:=\overleftarrow{F}^{(t)}\left(\left(3/4\right)^{t}n\right)-\overleftarrow{F}^{(t)}\left(\left(3/4\right)^{t}(n-1)\right)$
@@
mass at $n$; and by re-scaling by a factor of $(\frac{4}{3})^{t}$, any new mass has the same units over time. If there is an un-merging at the $n$-th index, then an additional $\overleftarrow{Z}^{(t)}_{n}$ is created, and no new mass is created otherwise. However, any new mass is also shrunken by a factor of $\frac{3}{4}$ at time-step $t+1$. Thus, the contribution of the mass at index $n$ at time $t+1$ is given by $\left(\frac{3}{4}\right)^{t+1}\overleftarrow{Z}^{(t)}_{n}(\mathds{1}_{n\in\overleftarrow{\rho}^{(t+1)}}-\frac{1}{3})$. The total additional mass at time $(t+1)$ contributed from indices contributing to $\overleftarrow{F}^{(t)}(x)$ is then given by summing over $n$.
@@
By a similar re-scaling argument, there are $|\overleftarrow{\rho}^{t+1}\cap[0,(\frac{4}{3})^{t}x]|$ un-mergings in this interval; and once again, those are scaled by an additional factor of $\frac{3}{4}$ at time $t+1$.

### v2

@@ 4 Proofs of Main Results
and thus $\overleftarrow{L}^{(t)}\to 0$ in $L^{2}$. It follows that $\langle\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)},\overleftarrow{\Gamma}^{(0)}\rangle\to\overleftarrow{M}^{(\infty)}$ in $L^{2}$. This establishes the one-dimensional distribution convergence since the choice of $\overleftarrow{\eta}^{(0)}=e_{0}$ is without loss of generality.
We now extend the argument to finite-dimensional distributions. Let $F\subset\mathds{Z}$ be finite and take
$\overleftarrow{\eta}^{(0)}=\sum_{i\in F}e_{i}.$
Using the same renewal environment $(\rho^{(t)})_{t\geqslant 1}$, the corresponding reverse-time weight processes form a finite collection of martingales, each of which satisfies the same a.s. and $L^{2}$ convergence bounds as in the case $\overleftarrow{\eta}^{(0)}=e_{0}$ (by shift-invariance of the dynamics and the previous argument). Therefore, any finite linear combination of the associated dual pairings converges in $L^{2}$, and hence in distribution. This yields convergence of all finite-dimensional marginals.
Finally, we show that the limiting gap distribution has exponential tails. Notice that the limiting gap distribution is given by $\overleftarrow{M}^{(\infty)}$, so it suffices to obtain a bound via the dynamics on $\overleftarrow{M}$. We will prove that for some $\lambda>0$, $\operatorname{\mathds{E}}e^{\lambda\overleftarrow{M}^{(\infty)}}<\infty$.
Note that the random variables $\overleftarrow{M}_{t}$ are bounded for all $t\in\{1,2,\ldots\}$ and thus have finite MGFs. Using the decomposition $\overleftarrow{M}_{t+1}-\overleftarrow{M}_{t}=\overleftarrow{N}^{(t+1)}$ from equation (4.1), the MGF bound from Lemma 3.10, and the bound from equation (4.2), we have for some constants $c,C>0$ and all $\left|\lambda\right|\leqslant c,t\geqslant 0$
$\operatorname{\mathds{E}}(e^{\lambda(\overleftarrow{M}_{t+1}-\overleftarrow{M}_{t})}\mid\mathcal{F}_{t})\leqslant\exp(C\cdot\lambda^{2}\cdot(3/4)^{t}\overleftarrow{M}_{t}).$
Thus, we have for $0<\lambda\leqslant c,t\geqslant 0$,
$\operatorname{\mathds{E}}e^{\lambda\overleftarrow{M}_{t+1}}\leqslant\operatorname{\mathds{E}}e^{\lambda(1+C(3/4)^{t}\lambda)\overleftarrow{M}_{t}}.$
Now fix $\lambda_{0}\in[0,c]$ and recursively define $\lambda_{t+1}$ as a function of $\lambda_{t}$ by
$\lambda_{t}=\lambda_{t+1}\cdot(1+C\lambda_{t+1}\cdot(3/4)^{t}),t\geqslant 0.$
We have
$\operatorname{\mathds{E}}e^{\lambda_{t+1}\overleftarrow{M}_{t+1}}\leqslant\operatorname{\mathds{E}}e^{\lambda_{t}\overleftarrow{M}_{t}}\text{ for all }t\geqslant 0.$
Clearly, $\lambda_{t}$ is decreasing and converges to a constant $\lambda_{\infty}\geqslant\lambda_{0}\prod_{t=0}^{\infty}(1+C\lambda_{0}(3/4)^{t})^{-1}>0$. By Fatou’s lemma, we obtain a finite MGF bound for $M_{\infty}$, completing the proof. $\blacksquare$
@@
The proof follows the same martingale argument as in Theorem 3.1, with only a modification of the reverse-time update rule to encode cluster-size growth rather than gap mass. Specifically, in the reverse-time construction, instead of using the tuples $(\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$ and $(\overleftarrow{\eta_{i}}^{(t)},2\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$, we instead use the tuples $(\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$ and $(0,2\overleftarrow{\eta_{i}}^{(t)},2\overleftarrow{\eta_{i}}^{(t)})$. All subsequent steps (martingale decomposition, $L^{2}$ bounds, and convergence) are unchanged, since they depend only on the same renewal-environment covariance estimates and the same scaling structure. The exponential tail decay also follows the same argument as used in the proof of Theorem 3.1. $\blacksquare$
@@
It suffices to verify the duality identity for initial conditions of the form $\overleftarrow{\eta}^{(0)}=e_{i}$, $i\in\mathds{Z}$, since both sides are linear in $\overleftarrow{\eta}^{(0)}$ (for finitely supported initial configurations), and the general statement follows by linearity of expectation. Recall that, for $h(\cdot,\cdot)$ defined as the inner product in the corollary statement, the Markov processes $\overrightarrow{\Gamma}^{(\cdot)}$ and $\overleftarrow{\eta}^{(\cdot)}$ are dual if for all $t$, $\overrightarrow{\Gamma}^{(0)}$, and $\overleftarrow{\eta}^{(0)}$, we have
$\operatorname{\mathds{E}}_{\overrightarrow{\Gamma}^{(0)}}h(\overrightarrow{\Gamma}^{(t)},\overleftarrow{\eta}^{(0)})=\operatorname{\mathds{E}}_{\overleftarrow{\eta}^{(0)}}h(\overrightarrow{\Gamma}^{(0)},\overleftarrow{\eta}^{(t)}).$
For $\overleftarrow{\eta}^{(0)}=e_{i}$, the result is shown in the proof of Theorem 3.1. $\blacksquare$
@@
This result follows by taking the proofs of Theorems 3.1 and 3.3, and coupling the point processes used to determine the un-merging indices. $\blacksquare$
4.5 Proof of Theorem 3.5
We prove convergence of $\overleftarrow{F}^{(t)}$ by first establishing control of the moment generating functions $G^{(t)}$, and then identifying the limiting distribution function through the recursive dynamics in Lemma 3.6. Indeed, define for $s\in\mathds{R}$,
$G^{(t)}(s):=\int_{\mathds{R}}e^{sx}d\overleftarrow{F}^{(t)}(x).$
Using Lemma 3.6, we get:
$G^{(t+1)}(s)$ $=\int_{\mathds{R}}e^{sx}d\overleftarrow{F}^{(t+1)}(x)$
$=\int_{\mathds{R}}e^{s(x+\beta^{(t+1)}_{x})}d\overleftarrow{F}^{(t+1)}(x+\beta^{(t+1)}_{x})$
$=\int_{\mathds{R}}e^{sx+s\beta^{(t+1)}_{x}}d\overleftarrow{F}^{(t)}(x)+\int_{\mathds{R}}e^{sx+s\beta^{(t+1)}_{x}}d\alpha^{(t+1)}_{x}$
$=:H_{\beta}^{(t+1)}(s)+H_{\alpha}^{(t+1)}(s),$
where we treat $\alpha^{(t)}_{x}$ as a signed measure. Here, the second equality follows because the mapping $x\mapsto x+\beta^{(t+1)}_{x}$ is monotone increasing in $x$: we have $x+\beta^{(t+1)}_{x}=\frac{2}{3}x+N_{\rho^{(t+1)}}(x)$, where $N_{\rho^{(t+1)}}(\cdot)$ accounts for any positive jumps due to $\rho^{(t+1)}$.
Observe that for $x\in\mathds{R}$,
$\alpha_{x}^{(t+1)}$ $=-\frac{1}{4}\overleftarrow{F}^{(t)}(x)+\frac{3}{4}\sum_{n\leqslant(4/3)^{t}x}\overleftarrow{f}^{(t)}((3/4)^{t}n)\cdot\mathds{1}_{n\in\rho^{(t+1)}},$
and that
$H_{\beta}^{(t+1)}(s)$ $:=\int_{\mathds{R}}e^{sx+s\beta_{x}^{(t+1)}}d\overleftarrow{F}^{(t)}(x)=\sum_{n\in\mathds{N}}\exp\left(\left(\frac{3}{4}\right)^{t}sn+s\beta^{(t+1)}_{(3/4)^{t}n}\right)\overleftarrow{f}^{(t)}\left(\left(\frac{3}{4}\right)^{t}n\right),$
and thus we can re-express:
$G^{(t+1)}(s)$ $=\frac{3}{4}\sum_{n\in\mathds{N}}\exp\left(\left(\frac{3}{4}\right)^{t}sn+s\beta^{(t+1)}_{(3/4)^{t}n}\right)\overleftarrow{f}^{(t)}\left(\left(\frac{3}{4}\right)^{t}n\right)$
$\qquad+\frac{3}{4}\sum_{n\in\mathds{N}}\exp\left(\left(\frac{3}{4}\right)^{t}sn+s\beta^{(t+1)}_{(3/4)^{t}n}\right)\overleftarrow{f}^{(t)}\left(\left(\frac{3}{4}\right)^{t}n\right)\mathds{1}_{n\in\rho^{(t+1)}}.$
For notational cleanliness, let $\hat{n}:=\left(3/4\right)^{t}n$. Conditioning on $\mathcal{F}_{t}$, the terms $\beta^{(t+1)}_{\hat{n}}$ are functions of the renewal process $\rho^{(t+1)}$ only, so Lemma 3.7 applies termwise to the exponential factors involving $\beta^{(t+1)}_{\hat{n}}$. Using Lemma 3.7, we have
@@
$\leqslant\frac{3}{4}\sum_{n\in\mathds{N}}e^{s\hat{n}}\overleftarrow{f}^{(t)}\left(\hat{n}\right)\left(1+\left(\frac{3}{4}\right)^{t}As\right)e^{As\hat{n}}$
$\qquad+\frac{3}{4}\sum_{n\in\mathds{N}}\frac{1}{3}e^{s\hat{n}}\overleftarrow{f}^{(t)}\left(\hat{n}\right)\left(1+\left(\frac{3}{4}\right)^{t}As\right)e^{As\hat{n}}$,
where the last inequality holds for all sufficiently small $s>0$. Finally, using Lemma 3.6, we get
$\operatorname{\mathds{E}}\bigl(G^{(t+1)}(s)\mid\mathcal{F}_{t}\bigr)\leqslant(1+A(3/4)^{t}s)\cdot G^{(t)}(s(1+A(3/4)^{t}s)).$ (4.3)
Now take $\theta_{0}\in(-1/A,\infty)$. Consider the recursion
$\theta_{t}=\theta_{t+1}(1+A(3/4)^{t}\theta_{t+1}),\qquad t\geqslant 0.$
Clearly, $\theta_{t}$ is decreasing in $t$ and is lower bounded by $\theta_{0}\prod_{i=0}^{\infty}(1+A(3/4)^{i}\theta_{0})^{-1}.$ Thus $\theta_{*}:=\lim_{t\to\infty}\theta_{t}$ exists. By continuity, the map $\theta_{0}\mapsto\theta_{*}$ is a bijection. By construction of $(\theta_{t})$ and the conditional estimate (4.3), the adapted process
$\bigl(\theta_{t}\,G^{(t)}(\theta_{t})\bigr)_{t\geqslant 0}$
is a nonnegative supermartingale. Hence, it converges almost surely to a finite limit, which we denote by
$G^{(\infty)}(\theta_{*}):=\lim_{t\to\infty}G^{(t)}(\theta_{t}).$
To identify the limit, we use the previously established convergence of the scaled gap distributions together with the a.s. control on total mass and support length encoded by the reverse-time construction. In particular, on an event of probability one, the family $\{\overleftarrow{F}^{(t)}\}_{t\geqslant 0}$ is tight and has uniformly bounded total mass and support along large times. Consequently, for any sequence $t_{k}\to\infty$, there exists a further subsequence (not relabeled) such that $\overleftarrow{F}^{(t_{k})}$ converges weakly to some random limit $\widetilde{F}^{(\infty)}$.
By Theorem 3.1, the limiting gap law is unique; therefore, the total mass of $\widetilde{F}^{(\infty)}$ must agree in distribution with that of $\overleftarrow{F}^{(\infty)}$, and the recursive construction forces the same identification for the full distribution function. Hence, every subsequential weak limit coincides (almost surely) with $\overleftarrow{F}^{(\infty)}$. It follows that the full sequence $\overleftarrow{F}^{(t)}$ converges weakly almost surely to $\overleftarrow{F}^{(\infty)}$. $\blacksquare$
5 Proofs of Technical Results
5.1 Proof of Lemma 3.6
@@ 5 Proofs of Technical Results
Fix $t$ and consider some fixed integer $n\in\left[0,\left(4/3\right)^{t}\right]$. There is mass $\overleftarrow{f}^{(t)}((3/4)^{t}n)$ at $n$. By rescaling by a factor of $(\frac{4}{3})^{t}$, any new mass has the same units over time. If there is an un-merging at the $n$-th index, then an additional $\overleftarrow{f}^{(t)}((3/4)^{t}n)$ is created, and no new mass is created otherwise. However, any new mass is also shrunk by a factor of $\frac{3}{4}$ at time step $t+1$. Thus, the contribution of the mass at index $n$ at time $(t+1)$ is given by $\left(\frac{3}{4}\right)^{t+1}\cdot\overleftarrow{f}^{(t)}((3/4)^{t}n)\cdot(\mathds{1}_{n\in\overleftarrow{\rho}^{(t+1)}}-1/3)$. The total additional mass at time $(t+1)$ contributed from indices contributing to $\overleftarrow{F}^{(t)}(x)$ is then given by summing over $n$.
@@
By a similar re-scaling argument, there are $|\overleftarrow{\rho}^{t+1}\cap[0,(4/3)^{t}x]|$ un-mergings in this interval, and once again, those are scaled by an additional factor of $\frac{3}{4}$ at time $t+1$.